OpenStudy (anonymous):

A 20 kg block of ice slides without friction down a long hill. The start of the hill is 7.5 m higher than the bottom of the hill, and the path traveled is 18.0 m. Find the speed of the block when it reaches the bottom of the hill. Im not really sure where to start on this one

OpenStudy (anonymous):



OpenStudy (anonymous):

OK so the formula i found is mgh=mv²/2 v=sqrt(2gh)

OpenStudy (anonymous):

2as = vf2 - vi2

OpenStudy (anonymous):

Yup try that.... and this also.... Vf = final velocity and Vi = initial velocity.= 0 and s = h \[2ah = v _{f}^{2} - (0)^{2}\]

OpenStudy (anonymous):

and a =g

OpenStudy (anonymous):

ok so with the sqrt(2gh) i got 12.1 which i think is right

OpenStudy (anonymous):

hahaha both answer will be same dude... :D

OpenStudy (anonymous):

Yay well atleast i got it right the other formula seems so confusing but that might be because its 3 am

OpenStudy (anonymous):

its not confusing, its 3rd equation of motion which i used. and that's right because we are taliking about initial and final velocity. but formula you used that is in terms of average velocity... well both ways are right.

OpenStudy (anonymous):

Lol may not be confusing to you but my 3 am self is way different from my normal 7-8 hours of sleep self but thanks

OpenStudy (anonymous):

:) welcome